# SwiftVideo: A Unified Framework for Few-Step Video Generation through Trajectory-Distribution Alignment

### v1

@@ Introduction
Table 2: Ablation studies. We set the sampling steps to 4. a) Effect of CCD. b) Effect of DA. c) Effect of TA. d) The time schedule.
Effect of Synthetic Preference Dataset in Trajectory Alignment
We compared two preference dataset strategies: synthetic preference datasets versus using real data as preferred samples with low-step generated data as less preferred samples. As shown in Figure 6, synthetic preference datasets enable rapid improvement in few-step inference quality (red line), while real data as preferred samples leads to slow convergence (blue line). This occurs because synthetic preferred data implicitly contains trajectory information, allowing the model to quickly enhance low-step generation by aligning with high-step inference trajectories.
Effect of Reflow Loss in Trajectory Alignment
We investigated the impact of reflow loss by comparing models trained using $\mathcal{L}_{\textnormal{dpo}}$ alone versus those trained with the combined objective $\mathcal{L}_{\textnormal{DPO}}+\lambda_{\textnormal{RF}}\cdot\mathcal{L}_{\textnormal{RF}}$. As shown in Figure 6 (red line vs. green line), incorporating reflow loss enhances peak performance and stabilizes training.
Conclusion
@@
Supplementary Material
In this supplementary material, we provide additional details to complement the main paper. It includes discussions with concurrent work MeanFlow (Geng et al. 2025), computational details of continuous-time consistency distillation, and supplementary information on trajectory alignment. Furthermore, we present quantitative comparison results on the VBench-I2V dataset and qualitative results on external data, demonstrating the strong generalization capability of our method. Additionally, we provide ablation studies that validate the robustness of hyperparameter tuning in trajectory alignment.
Discussion with MeanFlow
MeanFlow model (Geng et al. 2025) is a concurrent work that presents a one-step image generation model trained from scratch. As introduced in the Preliminary section, the velocity modeled in Flow Matching represents the instantaneous velocity. The marginal velocity field $\bm{v}(\mathbf{x}_{t},t)$ is defined as $\bm{v}(\mathbf{x}_{t},t)=\frac{d\mathbf{x}_{t}}{dt}$, where represents the time derivative. And in rectified flow $\bm{v}(\mathbf{x}_{t},t)=\mathbf{x}_{1}-\mathbf{x}_{0}$. Unlike the instantaneous velocity used in Flow Matching model, this work introduces the concept of average velocity to describe flow fields and leverages the identity between average and instantaneous velocities to train the one-step image generation model. In MeanFlow model, the average velocity $u$ is :
@@
Following the definition of average velocity in (Geng et al. 2025), $u(\mathbf{x}_{t},t)$ represents the average velocity between the time steps 0 and t, so we have $u(\mathbf{x}_{t},t)t=\mathbf{x}_{t}-\mathbf{x}_{0}$. Now we differentiate both sides with respect to t:
@@
Figure 7: FVD (Ge et al. 2024) evaluation of the distilled model using Euler solver inference on OpenVid-1M test set (Nan et al. 2025). The model shows significant quality improvements with increasing inference steps, particularly in the low-step regime ($\textnormal{steps}\leq 8$).
@@
Figure 8: Win Diff during training for DPO with and without reflow loss regularization. The DPO with reflow loss (red line) effectively stabilizes the Win Diff compared to standard DPO (blue line).
Details of Trajectory Alignment
@@
We analyzed the performance of distribution-aligned distilled models using Euler solver inference at different sampling steps. As shown in Figure 7, the distilled model exhibits significant quality improvements with increasing step counts in the low-step regime ($\textnormal{steps}\leq 8$). This empirical observation motivates our subsequent employment of the Direct Preference Optimization (DPO) algorithm (Rafailov et al. 2023) to align the model with high-step trajectories.
@@
Additionally, we analyzed the undesired training dynamics when employing DPO loss alone for trajectory alignment. To evaluate model performance relative to the reference model, we define the win difference (Win Diff) as $\|\mathbf{v}^{w}-\mathbf{v}_{\theta}(\mathbf{x}_{t}^{w},t)\|^{2}-\|\mathbf{v}^{w}-\mathbf{v}_{\mathrm{ref}}(\mathbf{x}_{t}^{w},t)\|^{2}$, which measures the difference in velocity prediction errors between the training model and reference model on preferred data. As shown in Figure 8 (blue line), training with only DPO loss leads to a continuous increase in Win Diff throughout training iterations, indicating that the model diverges from the intended trajectory and performs worse than the reference model on preferred samples. This occurs because DPO loss can decrease even when the model’s velocity prediction accuracy degrades more on dispreferred data than on preferred data, resulting in suboptimal trajectory alignment. To address this issue, we introduce an additional reflow loss on preferred data as regularization to preserve high prediction fidelity. As shown in Figure 8 (red line), incorporating the reflow loss as a regularization term effectively prevents the Win Diff from increasing during training, enabling the model to better align with preferred trajectories while diverging from dispreferred ones.
More Quantitative Results
@@
In addition to evaluation on the Openvid-1M test set (Nan et al. 2025), we conduct comparative analysis against other existing distillation methods on the VBench-I2V (Huang et al. 2024) dataset. Following the experimental setup in the main text, we generate videos at $832\times 480$ resolution and 24 FPS during inference. And all experiments in this paper employ a fixed CFG (Ho and Salimans 2022) scale of 3.0.
@@
As demonstrated in Table 3, our method significantly outperforms other existing state-of-the-art distillation methods across all quantitative metrics for both 4-step and 2-step inference. Specifically, regarding frame quality, our 4-step approach achieves results that closely approximate the teacher model’s 25-step performance, while our 2-step method even surpasses the 4-step results of LCM (Luo et al. 2023), PCM (Wang et al. 2024), and DMD2 (Yin et al. 2024a). Furthermore, our method exceeds the teacher model’s 25-step performance on image-conditioning scores. These results demonstrate that we achieve state-of-the-art performance on the VBench-I2V dataset, effectively showcasing the strong generalization capability of our approach.
@@
More Qualitative Results
@@
More Ablation Studies

### v2

@@ 5.3.2 Qualitative Comparisons
Table 2: Ablation studies. We set the sampling steps to 4. a) Effect of CCD and DA. b) Effect of TA. c) The time schedule.
5.3.3 Effect of Continuous-time Consistency Distillation (CCD)
To validate continuous-time consistency distillation effectiveness, we evaluate three training strategies: direct distribution alignment (DA), discrete-time consistency distillation (DCD) followed by DA, and continuous-time consistency distillation (CCD) followed by DA. Table 6(a) demonstrates that consistency distillation combined with distribution alignment achieves superior performance. DA alone produces excessive trajectory deviation from the original model. Furthermore, CCD outperforms DCD through more accurate ODE trajectory preservation.
5.3.4 Effect of Distribution Alignment (DA)
To validate the effectiveness of distribution alignment, we conduct ablation studies using only continuous-time consistency loss for distillation. As shown in Table 6(a), DA yields superior performance by directly approximating the real data distribution, which enhances visual fidelity under few-step sampling.
5.3.5 Effect of Trajectory Alignment (TA)
Comparing distilled models with and without trajectory alignment, Table 6(b) demonstrates that model refined through TA generated higher-quality videos. This improvement stems from using DPO to align the model with high-step inference trajectories, thereby enhancing low-step inference quality.
5.3.6 Time Samplers
Prior work (Esser et al., 2024) demonstrates that the distribution of timestep affects generation quality. We train models using CCD with DA under different timestep distributions. As shown in Table 6(c), the logit-normal timestep sampler achieves superior results.
5.3.7 Effect of Synthetic Preference Dataset in Trajectory Alignment
We compared two preference dataset strategies: synthetic preference datasets versus using real data as preferred samples with low-step generated data as less preferred samples. As shown in Figure 5(b), synthetic preference datasets enable rapid improvement in few-step inference quality (red line), while real data as preferred samples leads to slow convergence (blue line). This occurs because synthetic preferred data implicitly contains trajectory information, allowing the model to quickly enhance low-step generation by aligning with high-step inference trajectories.
5.3.8 Effect of Reflow Loss in Trajectory Alignment
We investigated the impact of reflow loss by comparing models trained using $\mathcal{L}_{\textnormal{dpo}}$ alone versus those trained with the combined objective $\mathcal{L}_{\textnormal{DPO}}+\lambda_{\textnormal{RF}}\cdot\mathcal{L}_{\textnormal{RF}}$. As shown in Figure 5(b) (red line vs. green line), incorporating reflow loss enhances peak performance and stabilizes training.
6 Conclusions
@@
Appendix A Discussion with MeanFlow
MeanFlow model [37] is a concurrent work that presents a one-step image generation model trained from scratch. As introduced in the Preliminary section, the velocity modeled in Flow Matching represents the instantaneous velocity. The marginal velocity field $\bm{v}(\mathbf{x}_{t},t)$ is defined as $\bm{v}(\mathbf{x}_{t},t)=\frac{d\mathbf{x}_{t}}{dt}$, where represents the time derivative. And in rectified flow $\bm{v}(\mathbf{x}_{t},t)=\mathbf{x}_{1}-\mathbf{x}_{0}$. Unlike the instantaneous velocity used in Flow Matching model, this work introduces the concept of average velocity to describe flow fields and leverages the identity between average and instantaneous velocities to train the one-step image generation model. In MeanFlow model, the average velocity $u$ is :
@@ Appendix A Discussion with MeanFlow
Following the definition of average velocity in [37], $u(\mathbf{x}_{t},t)$ represents the average velocity between the time steps 0 and t, so we have $u(\mathbf{x}_{t},t)t=\mathbf{x}_{t}-\mathbf{x}_{0}$. Now we differentiate both sides with respect to t:
@@ A.1 JVP Computation Detail
(a) FVD [33] evaluation of the distilled model using Euler solver inference on OpenVid-1M test set [8]. The model shows significant quality improvements with increasing inference steps, particularly in the low-step regime ($\textnormal{steps}\leq 8$).
@@
(b) Win Diff during training for DPO with and without reflow loss regularization. The DPO with reflow loss (red line) effectively stabilizes the Win Diff compared to standard DPO (blue line).
Appendix B Details of Trajectory Alignment
@@
We analyzed the performance of distribution-aligned distilled models using Euler solver inference at different sampling steps. As shown in Figure 7(a), the distilled model exhibits significant quality improvements with increasing step counts in the low-step regime ($\textnormal{steps}\leq 8$). This empirical observation motivates our subsequent employment of the Direct Preference Optimization (DPO) algorithm [9] to align the model with high-step trajectories.
@@
Additionally, we analyzed the undesired training dynamics when employing DPO loss alone for trajectory alignment. To evaluate model performance relative to the reference model, we define the win difference (Win Diff) as $\|\mathbf{v}^{w}-\mathbf{v}_{\theta}(\mathbf{x}_{t}^{w},t)\|^{2}-\|\mathbf{v}^{w}-\mathbf{v}_{\mathrm{ref}}(\mathbf{x}_{t}^{w},t)\|^{2}$, which measures the difference in velocity prediction errors between the training model and reference model on preferred data. As shown in Figure 7(b) (blue line), training with only DPO loss leads to a continuous increase in Win Diff throughout training iterations, indicating that the model diverges from the intended trajectory and performs worse than the reference model on preferred samples. This occurs because DPO loss can decrease even when the model’s velocity prediction accuracy degrades more on dispreferred data than on preferred data, resulting in suboptimal trajectory alignment. To address this issue, we introduce an additional reflow loss on preferred data as regularization to preserve high prediction fidelity. As shown in Figure 7(b) (red line), incorporating the reflow loss as a regularization term effectively prevents the Win Diff from increasing during training, enabling the model to better align with preferred trajectories while diverging from dispreferred ones.
Appendix C More Quantitative Results
@@
In addition to evaluation on the Openvid-1M test set [8], we conduct comparative analysis against other existing distillation methods on the VBench-I2V [34] dataset. Following the experimental setup in the main text, we generate videos at $832\times 480$ resolution and 24 FPS during inference. And all experiments in this paper employ a fixed CFG [38] scale of 3.0.
@@
As demonstrated in Table 3, our method significantly outperforms other existing state-of-the-art distillation methods across all quantitative metrics for both 4-step and 2-step inference. Specifically, regarding frame quality, our 4-step approach achieves results that closely approximate the teacher model’s 25-step performance, while our 2-step method even surpasses the 4-step results of LCM [3], PCM [4], and DMD2 [5]. Furthermore, our method exceeds the teacher model’s 25-step performance on image-conditioning scores. These results demonstrate that we achieve state-of-the-art performance on the VBench-I2V dataset, effectively showcasing the strong generalization capability of our approach.
@@
Appendix D More Qualitative Results
@@
Appendix E More Ablation Studies
E.1 Effect of $\lambda_{RF}$n Trajectory Alignment